# HOTFormerLoc: Hierarchical Octree Transformer for Versatile Lidar Place Recognition Across Ground and Aerial Views

### v1

@@ Abstract
We present HOTFormerLoc, a novel and versatile Hierarchical Octree-based TransFormer, for large-scale 3D place recognition in both ground-to-ground and ground-to-aerial scenarios across urban and forest environments. We propose an octree-based multi-scale attention mechanism that captures spatial and semantic features across granularities. To address the variable density of point distributions from spinning lidar, we present cylindrical octree attention windows to reflect the underlying distribution during attention. We introduce relay tokens to enable efficient global-local interactions and multi-scale representation learning at reduced computational cost. Our pyramid attentional pooling then synthesises a robust global descriptor for end-to-end place recognition in challenging environments. In addition, we introduce CS-Wild-Places, a novel 3D cross-source dataset featuring point cloud data from aerial and ground lidar scans captured in dense forests. Point clouds in CS-Wild-Places contain representational gaps and distinctive attributes such as varying point densities and noise patterns, making it a challenging benchmark for cross-view localisation in the wild. HOTFormerLoc achieves a top-1 average recall improvement of 5.5% – 11.5% on the CS-Wild-Places benchmark. Furthermore, it consistently outperforms SOTA 3D place recognition methods, with an average performance gain of 5.8% on well-established urban and forest datasets. The code and CS-Wild-Places benchmark is available at https://csiro-robotics.github.io/HOTFormerLoc.
@@ 3.2 Hierarchical Octree Transformer
The effect of this on the 2D-equivalent of an octree is seen in Fig. 3, which visualises attention windows for two concentric rings of points with differing densities (mimicking the behaviour of spinning lidar in outdoor scenes). Cartesian octree attention windows cover a uniform area and ignore the underlying point density, whereas the cylindrical octree attention windows respect the point distribution, with fine-grained windows near the sensor, and sparser windows at lower-density regions further from away. We demonstrate the effectiveness of this approach on point clouds captured in natural environments in Tab. 8. See Supplementary Material for further visualisations.
@@ 5 Experiments
PointNetVLAD [48] 35.6 41.5
TransLoc3D [56] 58.2 69.0
MinkLoc3Dv2 [28] 67.1 76.7
CrossLoc3D [13] 74.1 83.0
HOTFormerLoc (Ours) 80.9 95.2
Table 2: Comparison of SOTA on CS-Campus3D [13] with ground-only queries, and ground + aerial database.
@@ 5.1 Comparison with SOTA
CS-Campus3D: In Tab. 2, we present the evaluation results on CS-Campus3D, training our method for 300 epochs with a LR of $5e^{-4}$, reduced by a factor of 10 after 250 epochs. Our approach shows an improvement of $6.8\%$ and $12.2\%$ in AR@1 and AR@1%, respectively. Most notably, we exceed the performance of CrossLoc3D [13], which employs a diffusion-inspired refinement step to specifically address the cross-source challenge. This highlights the versatility of our hierarchical attention approach, which is capable of learning general representations to achieve SOTA performance in both single- and cross-source settings.
@@ 6 Conclusion and Future Work
Acknowledgements: We acknowledge support of the Terrestrial Ecosystem Research Network (TERN), supported by the National Collaborative Infrastructure Strategy (NCRIS). This work was partially funded through the CSIRO’s Digital Water and Landscapes initiative (3D-AGB project). We thank the Research Engineering Facility (REF) team at QUT for their expertise and research infrastructure support and Hexagon for providing SmartNet RTK corrections for precise surveying.
@@ 7.3 Pyramid Attentional Pooling
C = 256 M = 10 35.4 M 62 / 270 ms 92.1 ($\uparrow$2.1) 80.9 ($\uparrow$6.8) 60.5 ($\uparrow$8.5)
C = 256 M = 8 28.9 M 50 / 250 ms 91.8 ($\uparrow$1.8) 75.9 ($\uparrow$1.8) 58.9 ($\uparrow$6.9)
C = 256 M = 6 22.6 M 41 / 228 ms 91.5 ($\uparrow$1.5) 72.3 ($\downarrow$1.8) 57.6 ($\uparrow$5.6)
C = 192 M = 8 16.7 M 40 / 192 ms 90.8 ($\uparrow$0.8) 75.6 ($\uparrow$1.5) 58.1 ($\uparrow$6.1)
@@ 7.4 HOTFormerLoc Ablations
We provide ablations on the number of HOTFormer blocks and channel size in Tab. 10. HOTFormerLoc maintains SOTA performance with fewer parameters than the full-sized model, outperforming MinkLoc3Dv2 by $8.5\%$ on CS-Campus3D with 16.7M params. This parameter count is similar to existing transformer-based LPR methods [56, 21], whilst outperforming them by $17.4\%$ on CS-Campus3D. We also report the runtime on dense point clouds from CS-Wild-Places, and the sparse point clouds from CS-Campus3D, with HOTFormerLoc achieving $40-62~{}ms$ inference time when limited to 4096 points.
@@ 7.5 Limitations and Future Work
As mentioned in Tab. 6, the runtime of HOTFormerLoc can be further improved through parallelisation. While our design is best suited for parallel implementation, currently, the H-OSA layers for each pyramid level are computed in serial. To unlock the full potential of our network design for optimal runtime, these layers can be combined into a single operation. Furthermore, the octree implementation used in HOTFormerLoc can be parallelised to enable more efficient octree construction, leading to further runtime improvements.

### v2

@@ Abstract
We present HOTFormerLoc, a novel and versatile Hierarchical Octree-based TransFormer, for large-scale 3D place recognition in both ground-to-ground and ground-to-aerial scenarios across urban and forest environments. We propose an octree-based multi-scale attention mechanism that captures spatial and semantic features across granularities. To address the variable density of point distributions from spinning lidar, we present cylindrical octree attention windows to reflect the underlying distribution during attention. We introduce relay tokens to enable efficient global-local interactions and multi-scale representation learning at reduced computational cost. Our pyramid attentional pooling then synthesises a robust global descriptor for end-to-end place recognition in challenging environments. In addition, we introduce CS-Wild-Places, a novel 3D cross-source dataset featuring point cloud data from aerial and ground lidar scans captured in dense forests. Point clouds in CS-Wild-Places contain representational gaps and distinctive attributes such as varying point densities and noise patterns, making it a challenging benchmark for cross-view localisation in the wild. HOTFormerLoc achieves a top-1 average recall improvement of 5.5% – 11.5% on the CS-Wild-Places benchmark. Furthermore, it consistently outperforms SOTA 3D place recognition methods, with an average performance gain of 4.9% on well-established urban and forest datasets. The code and CS-Wild-Places benchmark is available at https://csiro-robotics.github.io/HOTFormerLoc.
@@ 3.2 Hierarchical Octree Transformer
The effect of this on the 2D-equivalent of an octree is seen in Fig. 3, which visualises attention windows for two concentric rings of points with differing densities (mimicking the behaviour of spinning lidar in outdoor scenes). Cartesian octree attention windows cover a uniform area and ignore the underlying point density, whereas the cylindrical octree attention windows respect the point distribution, with fine-grained windows near the sensor, and sparser windows at lower-density regions further from away. We demonstrate the effectiveness of this approach on point clouds captured in natural environments in Tab. 8. See Fig. 6 in the Supplementary for further detail.
@@ 5 Experiments
PointNetVLAD [48] 19.1 43.6
TransLoc3D [56] 43.0 80.6
MinkLoc3Dv2 [28] 52.5 83.5
CrossLoc3D [13] 70.7 85.7
HOTFormerLoc (Ours) 80.4 94.9
Table 2: Comparison of SOTA on CS-Campus3D [13] with ground-only queries, and aerial-only database.
@@ 5.1 Comparison with SOTA
CS-Campus3D: In Tab. 2, we present the evaluation results on CS-Campus3D, training our method for 300 epochs with a LR of $5e^{-4}$, reduced by a factor of 10 after 250 epochs. Our approach shows an improvement of $9.7\%$ and $9.2\%$ in AR@1 and AR@1%, respectively. Most notably, we exceed the performance of CrossLoc3D [13], which employs a diffusion-inspired refinement step to specifically address the cross-source challenge. This highlights the versatility of our hierarchical attention approach, which is capable of learning general representations to achieve SOTA performance in both single- and cross-source settings.
@@ 6 Conclusion and Future Work
Acknowledgements
We acknowledge support of the Terrestrial Ecosystem Research Network (TERN), supported by the National Collaborative Infrastructure Strategy (NCRIS). This work was partially funded through the CSIRO’s Digital Water and Landscapes initiative (3D-AGB project). We thank the Research Engineering Facility (REF) team at QUT for their expertise and research infrastructure support and Hexagon for providing SmartNet RTK corrections for precise surveying.
@@ 7.3 Pyramid Attentional Pooling
C = 256 M = 10 35.4 M 62 / 270 ms 92.1 ($\uparrow$2.1) 80.4 ($\uparrow$9.7) 60.5 ($\uparrow$8.5)
C = 256 M = 8 28.9 M 50 / 250 ms 91.8 ($\uparrow$1.8) 75.5 ($\uparrow$4.8) 58.9 ($\uparrow$6.9)
C = 256 M = 6 22.6 M 41 / 228 ms 91.5 ($\uparrow$1.5) 71.9 ($\uparrow$1.2) 57.6 ($\uparrow$5.6)
C = 192 M = 8 16.7 M 40 / 192 ms 90.8 ($\uparrow$0.8) 75.2 ($\uparrow$4.5) 58.1 ($\uparrow$6.1)
@@ 7.4 HOTFormerLoc Ablations
We provide ablations on the number of HOTFormer blocks and channel size in Tab. 10. HOTFormerLoc maintains SOTA performance with fewer parameters than the full-sized model, outperforming MinkLoc3Dv2 by $22.7\%$ on CS-Campus3D and $6.1\%$ on CS-Wild-Places with just 16.7M params. This parameter count is similar to existing transformer-based LPR methods [56, 21], whilst outperforming them by $32.2\%$ on CS-Campus3D. We also report the runtime on dense point clouds from CS-Wild-Places, and the sparse point clouds from CS-Campus3D, with HOTFormerLoc achieving $40-62~{}ms$ inference time when limited to 4096 points.
@@ 7.5 Limitations and Future Work
As mentioned in Tab. 6, the runtime of HOTFormerLoc can be improved through parallelisation. While our design is best suited for parallel implementation, currently, the H-OSA layers for each pyramid level are computed in serial. To unlock the full potential of our network design for optimal runtime, these layers can be combined into a single operation. Furthermore, the octree implementation used in HOTFormerLoc can be parallelised to enable more efficient octree construction.